OpenStudy (anonymous):

3x+4y=-24 Find x and y intercept

OpenStudy (anonymous):

For x intercept put y=0 For y intercept put x=0

OpenStudy (anonymous):

or u can write the eq in intercept form (x/a) +(y/b)=1 here it is (3x/-24)+(4y/-24) =1 or (x/(-8))+ (y/(-6))=1 thus x-intercept -8 and y-intercpt -6

OpenStudy (anonymous):

so you just double it?

OpenStudy (anonymous):

Nope...It Does nt Mean...that for this kind of question..)